Vector A⃗ =4i^+3j^ and vector B⃗ =5i^−4j^+3k^.

What is the cross product A⃗ ×B⃗ ? Find the x-component.

Find the y-component.

Find the z-component.
Answer #1

Answer Given that A = un + 3 B=51-45 + 3 % cros product can be woutten os à B = (A) (B) sinon ÅXB - I i § ŕ 4 3 0 15 -4 3 ÀXB